# PolSAM: Polarimetric Scattering Mechanism Informed Segment Anything Model

### v1

@@ IV-A Revisiting SAM
Building on the success of SAM in the segmentation field, we adopt its image encoder and mask decoder as the backbone of our segmentation network, keeping the encoder frozen. Inspired by SAM’s prompt-based mechanism, which relies on geometric prompts such as points, boxes, or masks, we propose a novel approach that autonomously generates semantic prompts. By incorporating domain-specific cues, our method produces both sparse and dense semantic prompts, enabling the model to effectively leverage task-specific and data-specific information for enhanced segmentation performance. The detailed structure is described below.
IV-B PolSAM Overview
The overview of our proposed PolSAM is illustrated in Fig. 6. We adopt the SAM image encoder ($f_{\text{Enc}}$) and mask decoder ($f_{\text{Dec}}$), keeping their original structures unchanged. Pseudo-colored images offer rich texture details and strong visual representation, while MVD captures scattering mechanisms and conveys semantic information. To effectively exploit and integrate these data-specific features, we propose the Feature-Level Fusion Prompt (FFP) module, placed before the image encoder to fuse features from both inputs for a more comprehensive representation. Additionally, to address data modality incompatibility and better adapt the encoder to PolSAR data, we introduce trainable adapter modules into each layer of the frozen image encoder, enhancing feature representation and segmentation performance.
To provide effective semantic prompts to the decoder, we designed a progressive Semantic-Level Fusion Prompt (SFP) module. Each level targets different aspects of key information, enabling nuanced semantic integration. The process is described by the following equations:
@@ IV-B PolSAM Overview
IV-C Feature-Level Fusion Prompt
The design of FFP follows the concept proposed in [50], where a lightweight module with convolutions is inserted into the beginning of the image encoder. In this context, the pseudo-colored image and MVD image are denoted as $I_{1}$ and $I_{2}$, respectively, with their corresponding patch embeddings represented by $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$. Given the distinct characteristics of these two inputs, $\mathbf{z}_{1}$ is closely related to local details and texture in the image, while $\mathbf{z}_{2}$ reflects the scattering mechanism, which is more related to semantic information. Therefore, enhancing the spatial resolution of $\mathbf{z}_{1}$ is essential to emphasize local details, whereas $\mathbf{z}_{2}$ undergoes direct dimensionality reduction to retain its global semantic features.
@@ IV-C Feature-Level Fusion Prompt
Specifically, as shown in Fig. 7, FFP first takes $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ as input and performs dimension reduction using 1$\times$1 convolutions to obtain the latent features for each input. The latent features of $\mathbf{z}_{1}$ are then processed with a Softmax activation across spatial dimensions to emphasize important regions, followed by a multiplication operation to enhance those areas. Subsequently, the fusion features are projected back to their original dimension using 1$\times$1 convolutions, yielding the initial fusion prompt at the feature level. Here, we use $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$ to represent the three convolution layers. Therefore, the fusion prompt can be derived as:
@@
IV-D Adapter-Equipped Image Encoder
The previous $\mathbf{P}_{\mathrm{F}}$ is then added by $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ to input to the pre-trained image encoder. The encoder backbone is frozen during training and we apply the adapter technology proposed in [34] within each layer of the image encoder to adapt the feature extraction. In this way, the fusion prompt and the adapters are learned to satisfy the complementarity of different inputs by fine-tuning only a few parameters. Thus, we can detail the acquisition of $\mathbf{F}_{\text{fused}}$ as presented in Equation (2):
@@ IV-D Adapter-Equipped Image Encoder
IV-E Semantic-Level Fusion Prompt
Inspired by SAM’s promptable mechanism, we propose a novel SFP module. This module automatically generates sparse and dense prompt embeddings to provide semantic guidance for segmentation. The SFP module features a two-level progressive structure aimed at optimizing feature integration and enhancing semantic representation. This design ensures a balanced extraction of unique and complementary information from different inputs, leading to a refined understanding that enriches the segmentation process. By progressively refining feature interactions, the SFP module significantly improves the model’s ability to generate precise and contextually aware segmentation results.
First, the pseudo-colored images $I_{1}$ are processed using a basic convolutional block and two pooling layers to obtain the feature embedding $\mathbf{F}_{1}$, which is represented by $f_{\text{FE}}^{1}(\cdot)$, as shown in Fig. 6. Considering the strong correlation between MVD and semantic labels, for MVD, we directly adopt the CNN network from the SAM prompt encoder used for processing mask prompts. This layer sequentially downsamples the input through three convolution layers, with intermediate normalization and activation functions, reducing the input channels and mapping the final output to the target embedding dimension. To ensure that the semantic prompts are not affected by color encoding, we directly use the one-hot form of the MVD, denoted as $I^{\prime}_{2}$. We then obtain the feature embedding $\mathbf{F}_{2}$, represented by $f_{\text{FE}}^{2}(\cdot)$. The features $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ have the same dimension as $\mathbf{F}_{\text{fused}}$. The above can be denoted as:
@@ IV-E Semantic-Level Fusion Prompt
The proposed SFP is a progressive process, with two levels implemented using a shared module. To clarify, we denote them as SFP-1 and SFP-2. As shown in Fig. 6, each level of the SFP module receives three inputs, denoted as $f_{\text{SFP}}(x_{1},x_{2},x_{3})$. The inputs of the SFP-1 are the single input embedding $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$, together with the fused feature from encoder $\mathbf{F}_{\mathrm{fused}}$. It outputs the initial sparse and dense prompt $\mathbf{P}_{\mathrm{S,D}}^{(1)}$. Then, to make the semantic prompts more specific, SFP-2 combines the single-input features and the fusion features in an additive manner, and interacts with the semantic features $\mathbf{F}_{\mathrm{att}}$ after the mask decoder attention layer to generate the final sparse and dense prompt $\mathbf{P}_{\mathrm{S,D}}^{(2)}$, $\mathbf{P}_{\mathrm{S}}\in\mathbb{R}^{N\times C}$ and $\mathbf{P}_{\mathrm{D}}\in\mathbb{R}^{H\times W\times C}$. The process can be denoted as:
@@
The details of SFP is shown in Fig. 8. Given three arbitrary inputs of $x_{1},x_{2},x_{3}\in\mathbb{R}^{C\times H\times W}$, the channel feature split is applied to obtain two feature groups with identical dimensions of $\mathbb{R}^{C\times H\times W}$, denoted as $[T_{1}^{m},T_{2}^{m},T_{3}^{m}]$, $[T_{1}^{n},T_{2}^{n},T_{3}^{n}]$, respectively. It is realized by a linear layer to increase the dimension from $C$ to $2C$. By exchanging the query and value, two groups of enhanced features can be obtained with cross-attention modules, denoted as $[v_{1},v_{2}]$ and $[u_{1},u_{2}]$, with the dimension of $HW\times C$. Then, the sparse prompt and dense prompt are generated separately with the following operations:
$\mathbf{P}_{\mathrm{S}}$ $=\text{Norm}(\text{Linear}_{1}(0.5(v_{1}+u_{1})+\text{DimTrans}(x_{1})))$ (7)
@@
From the above descriptions, we can find that SFP module aims to fuse different inputs interactively so that the information is enriched to a certain extent. The two cross-attention modules consider different inputs as the query to obtain the attention weight for feature enhancement. In the SFP-1, $\mathbf{F}_{1}$, $\mathbf{F}_{2}$, and $\mathbf{F}_{\mathrm{fused}}$ are corresponding to $x_{1}$, $x_{2}$, and $x_{3}$. To this end, $u_{1}$ and $u_{2}$ demonstrate that the fusion features for segmentation($\mathbf{F}_{\mathrm{fused}}$) is enhanced by the single-input features, while $v_{1}$ and $v_{2}$ indicate the remarkable single-input features concerned by fusion information. The initial sparse and dense semantic prompts pass through the cross-attention module in the mask decoder to output $\mathbf{F}_{\mathrm{att}}$, which represents high-level semantic features. To this end, in the SFP-2, $u_{1}$, $u_{2}$ and $v_{1}$, $v_{2}$ indicates the result of the mutual enhancement between the fusion information and high-level semantic features.
The significance of the SFP module lies in its ability to enhance feature representation through interactive fusion, progressively refining semantic prompts. This ensures that critical details from each input are preserved while enriching the overall feature space. Furthermore, the mutual enhancement between fused information and high-level semantic features further refines the model’s understanding of feature semantics, leading to improved segmentation results.
TABLE I: Comparison of PolSAM performance on the PhySAR-Seg-1 dataset: methods above the dashed line are SAM-based, and those below are multimodal. Metrics include per-category IoU and overall performance. Best results are in bold, second-best underlined, and complete PolSAM is marked in red
Models IoU Per Category (%) Overall Metrics (%)
HQ-SAM[36] 22.05 4.92 10.76 1.05 1.01 27.34 31.49 25.70 11.19
SAM-LST[35] 32.26 60.20 79.60 57.80 54.73 66.14 72.72 72.70 58.45
Personalize-SAM[32] 44.65 65.12 83.99 60.93 73.29 69.8 78.26 79.39 66.30
Mobile-SAM[33] 22.91 52.09 77.28 42.85 44.24 55.62 64.18 64.84 49.17
RSAM-Seg[41] 49.55 69.32 85.63 66.15 74.60 73.23 81.18 81.67 69.75
PolSAM(w/o MVD) 51.22 69.64 86.45 68.98 78.52 76.21 82.62 83.37 71.84
[…]
SFAF-MA[52] 54.29 73.03 87.56 69.41 77.07 76.99 83.67 84.33 73.06
GMNet[53] 45.41 65.11 50.03 53.05 65.16 70.74 75.96 75.62 58.25
CMX[54] 36.52 65.07 46.87 63.58 64.5 60.18 69.15 73.24 56.12
LSNet[55] 28.87 59.83 79.45 53.69 47.04 57.07 67.55 72.76 54.49
PolSAM(w/ MVD) 55.44 73.97 87.97 69.62 78.66 75.65 83.70 84.71 73.55
TABLE II: Comparison of PolSAM performance on the PhySAR-Seg-2 dataset
Models IoU Per Category (%) Overall Metrics (%)
HQ-SAM[36] 7.91 15.89 0.77 0.11 0.00 6.88 7.98 44.71 29.69 21.88 10.53
SAM-LST[35] 29.22 78.26 18.27 15.55 27.44 32.67 47.75 93.38 74.19 58.29 42.82
Personalize-SAM[32] 28.61 65.63 14.68 5.74 1.74 32.06 47.06 94.62 70.24 50.39 36.27
Mobile-SAM[33] 28.86 68.54 13.18 5.91 1.98 33.98 50.74 93.87 70.76 49.07 37.13
RSAM-Seg[41] 28.11 73.45 12.47 11.24 17.99 33.32 35.64 94.58 68.59 56.38 38.35
[…]
SFAF-MA[52] 32.78 81.39 17.74 12.39 5.82 33.28 52.62 94.54 74.49 52.69 41.32
GMNet[53] 31.75 82.73 24.75 16.17 11.08 36.39 55.23 95.22 77.31 60.92 44.17
CMX[54] 29.21 73.12 14.20 22.02 2.14 36.93 41.32 94.90 69.75 53.84 39.23
LSNet[55] 27.70 75.36 18.31 0.12 19.93 29.06 46.14 94.06 70.68 53.11 38.83
IV-F Loss Function
@@
For PhySAR-Seg-2 dataset, where the class distribution is more unbalanced, we adopt the focal loss function to calculate the loss. The weight for each category is the inverse of its proportion in the dataset, enhancing the model’s focus on harder-to-classify, less frequent categories. The focal loss is given by:
@@
V Experiments and Discussions
This section proceeds as follows: a description of the experimental configuration, results analysis and discussion, including metric comparisons and the interpretation of visualization outcomes. Subsequent detailed analyses are presented, encompassing the ablation studies on module design, an evaluation of the effectiveness of semantic prompts along with a hyperparameter analysis of sparse semantic prompts, and discussions on the effectiveness of MVD.
V-A Experimental Settings
Implementation Details. The pre-trained ViT-B model in SAM is applied as our encoder backbone. And we load the pre-trained parameters of prompt encoder and mask decoder. The framework is trained with AdamW optimizer, where the hyperparameters of $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.999, respectively. The batch size and the epoch number are set to 12 and 350, respectively. The starting learning rate is set to 0.001. Warming-up is applied at the beginning of the 4% training process, and then the learning rate undergoes exponential decay with a factor of 0.9 as the number of epochs increases. The experiments are carried out with a single NVIDIA 4090Ti GPU.
V-B Comparison With State-of-the-art Methods
The results of PolSAM are compared with ten state-of-the-art (SOTA) image segmentation methods for each dataset. We use only the pseudo-colored images as input, denoted as PolSAM (w/o MVD), in comparison with five SAM-based methods, including those using optical [36, 35, 33, 32] and remote sensing images [41], since the original SAM is designed to accept only a single input. In addition, as our method utilizes two inputs, we also compare it with five multimodal fusion algorithms [51, 52, 53, 54, 55]. The comparison results are reported in Table I and Table II. The following is a detailed description of the results of the experiment.
Evaluation Metrics. For PhySAR-Seg-1 dataset, as shown in Table I, PolSAM outperforms other advanced SAM-based methods designed for both natural images and remote sensing data. Specifically, our method PolSAM(w/o MVD) surpasses the best-performing method, RSAM-Seg [41], with improvements of 1.44% in mAcc, 1.7% in mF1 score, and 2.09% in mIoU. Furthermore, when compared with other multimodal fusion approaches, PolSAM demonstrates significant superiority, achieving SOTA performance.
For PhySAR-Seg-2 dataset, as shown in Table II, the SAM-LST method [35] achieves the best performance among the SAM-based approaches, with an mIoU of 42.82%. In comparison, the proposed PolSAM (w/o MVD) achieves 45.30%, showing an improvement of 2.48%. Since the SAR data is significantly different from natural and optical remote sensing images, existing SAM-based models struggle to achieve optimal performance. When compared to multimodal fusion methods, PolSAM achieves a 0.48% higher mF1 score and a 2.38% improvement in mIoU over the GMNet approach [53], demonstrating its effectiveness in handling diverse data inputs.
Additionally, the metrics for each category of the two datasets are visualized in Fig. 9, providing a clear comparison of the results. The blue bars represent our PolSAM(w/o MVD) method, corresponding to the blue-themed SAM-based approaches, while the red bars indicate our PolSAM(w/ MVD) method, aligned with the red-themed multimodal methods. Our methods demonstrate three key strengths. First, both PolSAM (w/ MVD) and PolSAM (w/o MVD) consistently outperform other methods across most categories. Second, PolSAM (w/ MVD) generally delivers better overall performance than PolSAM (w/o MVD), highlighting the advantage of incorporating MVD. Finally, for the PhySAR-Seg-1 dataset, our method achieves the best results across nearly every category. For the PhySAR-Seg-2 dataset, particularly in the urban class where double-bounce scattering is dominant but the class proportion is small, our method demonstrates significant improvements compared to other approaches. This indicates that, even in challenging scenarios, effectively leveraging MVD information can substantially enhance segmentation accuracy.
@@ V-B Comparison With State-of-the-art Methods
Figure 10: Visualization of segmentation results on the PhySAR-Seg-1 dataset: each row displays a randomly selected image from the test set, while each column shows results from different models.
Figure 11: Visualization of segmentation results on the PhySAR-Seg-2 dataset: each row displays a randomly selected image from the test set, while each column shows results from different models.
Visualization. Fig. 10 showcases the visualization of segmentation outcomes for the PhySAR-Seg-1 dataset, comparing SAM-based and multimodal fusion-based methods. Similarly, Fig. 11 presents the segmentation result visualizations for the PhySAR-Seg-2 dataset. It can be seen that PolSAM has obtained the most refined segmentation results among those methods. Note that the MVD image provides distinct physical properties of terrain, which depict high-level semantic information better than pseudo-colored images. On the contrary, the pseudo-colored image contains richer texture information than MVD. The multi-input complementarities in feature and semantic level are clearly reflected. The proposed PolSAM can explicitly leverage the semantic information with the fusion prompt, which is used to hint the decoder for better segmentation.

### v2

@@ 4.1 Revisiting SAM
Building on SAM’s success in segmentation, we adopt its image encoder and mask decoder as our network backbone while keeping the encoder frozen. Inspired by SAM’s prompt-based mechanism that utilizes geometric prompts such as points, boxes, or masks, we propose a novel approach for autonomous semantic prompt generation. By incorporating domain-specific cues, our method produces both sparse and dense semantic prompts, enabling effective leverage of task-specific and data-specific information for enhanced segmentation performance. The detailed architecture is described below.
4.2 PolSAM Overview
Fig. 6 illustrates the overview of our proposed PolSAM. We adopt SAM’s image encoder ($f_{\text{Enc}}$) and mask decoder ($f_{\text{Dec}}$) while preserving their original structures. Pseudo-colored images provide rich texture details and strong visual representation, whereas MVD captures scattering mechanisms and conveys semantic information. To effectively exploit and integrate these data-specific features, we propose the Feature-Level Fusion Prompt (FFP) module, positioned before the image encoder to fuse features from both inputs for comprehensive representation. Additionally, to address data modality incompatibility and enhance encoder adaptation to PolSAR data, we introduce trainable adapter modules into each layer of the frozen image encoder, improving feature representation and segmentation performance.
To provide effective semantic prompts to the decoder, we design a progressive Semantic-Level Fusion Prompt (SFP) module. Each level targets distinct aspects of key information, enabling nuanced semantic integration. The process is described by the following equations:
@@ 4.2 PolSAM Overview
4.3 Feature-Level Fusion Prompt
The FFP design follows the concept proposed in Jia et al. (2022), where a lightweight convolutional module is inserted at the image encoder’s beginning. Here, the pseudo-colored image and MVD image are denoted as $I_{1}$ and $I_{2}$ respectively, with corresponding patch embeddings $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$. Given these inputs’ distinct characteristics, $\mathbf{z}_{1}$ relates closely to local details and image texture, while $\mathbf{z}_{2}$ reflects scattering mechanisms more associated with semantic information. Therefore, enhancing $\mathbf{z}_{1}$’s spatial resolution is essential for emphasizing local details, whereas $\mathbf{z}_{2}$ undergoes direct dimensionality reduction to preserve global semantic features.
As illustrated in Fig. 7, FFP takes $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ as input and performs dimension reduction using 1×1 convolutions to obtain latent features for each input. The latent features of $\mathbf{z}_{1}$ are then processed with Softmax activation across spatial dimensions to emphasize important regions, followed by multiplication to enhance these areas. Subsequently, the fused features are projected back to their original dimension using 1×1 convolutions, yielding the initial fusion prompt at the feature level. Here, $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$ represent the three convolution layers. The fusion prompt can thus be derived as:
@@ 4.3 Feature-Level Fusion Prompt
4.4 Adapter-Equipped Image Encoder
The resulting $\mathbf{P}_{\mathrm{F}}$ is then combined with $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ to form the input for the pre-trained image encoder. The encoder backbone remains frozen during training, while we apply the adapter technology proposed in Wu et al. (2023) within each encoder layer to adapt feature extraction. This approach enables the fusion prompt and adapters to learn complementarity between different inputs through fine-tuning of only a few parameters. The acquisition of $\mathbf{F}_{\text{fused}}$ can thus be detailed as presented in Equation (2):
@@ 4.4 Adapter-Equipped Image Encoder
4.5 Semantic-Level Fusion Prompt
Inspired by SAM’s promptable mechanism, we propose the SFP module that generates sparse and dense prompt embeddings to guide segmentation. Featuring a two-level progressive structure, SFP optimizes feature integration and enhances semantic representation. This design balances extraction of unique and complementary information from different inputs, enriching the segmentation process. Through progressive feature interaction refinement, the SFP module significantly improves segmentation accuracy and contextual awareness.
First, pseudo-colored images $I_{1}$ are processed using a basic convolutional block and two pooling layers to obtain feature embedding $\mathbf{F}_{1}$, denoted as $f_{\text{FE}}^{1}(\cdot)$, as shown in Fig. 6. Given the strong correlation between MVD and semantic labels, we directly adopt the CNN network from SAM’s prompt encoder for processing mask prompts. This network sequentially downsamples input through three convolution layers with intermediate normalization and activation functions, reducing input channels and mapping final output to the target embedding dimension. To ensure semantic prompts remain unaffected by color encoding, we directly use the one-hot form of MVD, denoted as $I^{\prime}_{2}$, to obtain feature embedding $\mathbf{F}_{2}$, represented by $f_{\text{FE}}^{2}(\cdot)$. Features $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ share the same dimension as $\mathbf{F}_{\text{fused}}$. This process can be denoted as:
@@ 4.5 Semantic-Level Fusion Prompt
The proposed SFP follows a progressive process with two levels implemented using a shared module. For clarity, we denote them as SFP-1 and SFP-2. As shown in Fig. 6, each SFP level receives three inputs, denoted as $f_{\text{SFP}}(x_{1},x_{2},x_{3})$. SFP-1 takes individual input embeddings $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$, along with the fused encoder feature $\mathbf{F}_{\text{fused}}$, to output initial sparse and dense prompts $\mathbf{P}_{\mathrm{S,D}}^{(1)}$. To generate more specific semantic prompts, SFP-2 additively combines single-input features with fusion features and interacts with semantic features $\mathbf{F}_{\text{att}}$ from the mask decoder attention layer to produce final sparse and dense prompts $\mathbf{P}_{\mathrm{S,D}}^{(2)}$, where $\mathbf{P}_{\mathrm{S}}\in\mathbb{R}^{N\times C}$ and $\mathbf{P}_{\mathrm{D}}\in\mathbb{R}^{H\times W\times C}$. This process can be denoted as:
@@
Fig. 8 shows the SFP details. Given three arbitrary inputs $x_{1},x_{2},x_{3}\in\mathbb{R}^{C\times H\times W}$, channel feature split is applied to obtain two feature groups with identical dimensions $\mathbb{R}^{C\times H\times W}$, denoted as $T_{1}^{m},T_{2}^{m},T_{3}^{m}]$ and $T_{1}^{n},T_{2}^{n},T_{3}^{n}]$ respectively. This is realized through a linear layer that increases dimension from $C$ to $2C$. By exchanging query and value, two groups of enhanced features are obtained using cross-attention modules, denoted as $v_{1},v_{2}]$ and $u_{1},u_{2}]$, with dimension $HW\times C$. The sparse and dense prompts are then generated separately through the following operations:
[…]
@@
From the above descriptions, the SFP module aims to fuse different inputs interactively, enriching information to a considerable extent. The two cross-attention modules treat different inputs as queries to obtain attention weights for feature enhancement. In SFP-1, $\mathbf{F}_{1}$, $\mathbf{F}_{2}$, and $\mathbf{F}_{\text{fused}}$ correspond to $x_{1}$, $x_{2}$, and $x_{3}$ respectively. Consequently, $u_{1}$ and $u_{2}$ demonstrate that fusion features for segmentation ($\mathbf{F}_{\text{fused}}$) are enhanced by single-input features, while $v_{1}$ and $v_{2}$ indicate remarkable single-input features attended by fusion information. The initial sparse and dense semantic prompts pass through the cross-attention module in the mask decoder to output $\mathbf{F}_{\text{att}}$, representing high-level semantic features. In SFP-2, $u_{1}$, $u_{2}$ and $v_{1}$, $v_{2}$ indicate results of mutual enhancement between fusion information and high-level semantic features.
The SFP module enhances feature representation through interactive input fusion and progressive semantic prompt refinement. This approach preserves critical details while enriching the feature space. Furthermore, mutual enhancement between fused information and high-level semantic features improves model understanding, leading to superior segmentation results.
4.6 Loss Function
@@
For the PhySAR-Seg-2 dataset, where class distribution is more imbalanced, we adopt focal loss to calculate the loss function. The weight for each category is the inverse of its dataset proportion, enhancing the model’s focus on harder-to-classify, less frequent categories. The focal loss is given by:
@@
5 Experiments and Discussions
This section presents experimental configuration, results analysis, and discussions, including metric comparisons and visualization interpretations. It encompasses ablation studies on module design, evaluation of semantic prompt effectiveness, hyperparameter analysis of sparse prompts, and discussion on MVD effectiveness.
5.1 Experimental Settings
Implementation Details. The pre-trained ViT-B model in SAM serves as our encoder backbone. The framework is trained using the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We set the batch size to 12, epochs to 350, and the initial learning rate to 0.001. A 4% training warm-up is applied, followed by exponential decay of the learning rate with a factor of 0.9. Experiments are conducted on a single NVIDIA 4090Ti GPU.
Table 1: Comparison of PolSAM performance on the PhySAR-Seg-1 dataset: methods above the dashed line are SAM-based, while those below are multimodal. Metrics include per-category IoU and overall performance. Best results are highlighted in bold, second-best are underlined, and complete PolSAM is marked in red.
Models IoU Per Category (%) Overall Metrics (%)
HQ-SAMKe et al. (2024) 22.05 4.92 10.76 1.05 1.01 27.34 31.49 25.70 11.19
SAM-LSTChai et al. (2023) 32.26 60.20 79.60 57.80 54.73 66.14 72.72 72.70 58.45
Personalize-SAMZhang et al. (2023e) 44.65 65.12 83.99 60.93 73.29 69.8 78.26 79.39 66.30
Mobile-SAMZhang et al. (2023a) 22.91 52.09 77.28 42.85 44.24 55.62 64.18 64.84 49.17
RSAM-SegZhang et al. (2024a) 49.55 69.32 85.63 66.15 74.60 73.23 81.18 81.67 69.75
PolSAM(w/o MVD) 51.22 69.64 86.45 68.98 78.52 76.21 82.62 83.37 71.84
MMSFormerReza et al. (2024) 46.19 68.79 85.14 63.35 67.57 72.64 79.52 80.03 67.28
SFAF-MAHe et al. (2023) 54.29 73.03 87.56 69.41 77.07 76.99 83.67 84.33 73.06
GMNetZhou et al. (2021) 45.41 65.11 50.03 53.05 65.16 70.74 75.96 75.62 58.25
CMXZhang et al. (2023c) 36.52 65.07 46.87 63.58 64.5 60.18 69.15 73.24 56.12
LSNetZhou et al. (2023) 28.87 59.83 79.45 53.69 47.04 57.07 67.55 72.76 54.49
PolSAM(w/ MVD) 55.44 73.97 87.97 69.62 78.66 75.65 83.70 84.71 73.55
Table 2: Comparison of PolSAM performance on the PhySAR-Seg-2 dataset.
Models IoU Per Category (%) Overall Metrics (%)
HQ-SAMKe et al. (2024) 7.91 15.89 0.77 0.11 0.00 6.88 7.98 44.71 29.69 21.88 10.53
SAM-LSTChai et al. (2023) 29.22 78.26 18.27 15.55 27.44 32.67 47.75 93.38 74.19 58.29 42.82
Personalize-SAMZhang et al. (2023e) 28.61 65.63 14.68 5.74 1.74 32.06 47.06 94.62 70.24 50.39 36.27
Mobile-SAMZhang et al. (2023a) 28.86 68.54 13.18 5.91 1.98 33.98 50.74 93.87 70.76 49.07 37.13
RSAM-SegZhang et al. (2024a) 28.11 73.45 12.47 11.24 17.99 33.32 35.64 94.58 68.59 56.38 38.35
MMSFormerReza et al. (2024) 30.84 72.00 14.28 8.60 0.00 37.93 44.60 93.80 72.24 51.51 37.75
SFAF-MAHe et al. (2023) 32.78 81.39 17.74 12.39 5.82 33.28 52.62 94.54 74.49 52.69 41.32
GMNetZhou et al. (2021) 31.75 82.73 24.75 16.17 11.08 36.39 55.23 95.22 77.31 60.92 44.17
CMXZhang et al. (2023c) 29.21 73.12 14.20 22.02 2.14 36.93 41.32 94.90 69.75 53.84 39.23
LSNetZhou et al. (2023) 27.70 75.36 18.31 0.12 19.93 29.06 46.14 94.06 70.68 53.11 38.83
@@ 5.1 Experimental Settings
Figure 10: Visualization of segmentation results on the PhySAR-Seg-1 (rows 1-3) and PhySAR-Seg-2 (rows 4-6) datasets: each row displays a randomly selected image from the test set, while each column shows results from different models.
5.2 Comparison With State-of-the-art Methods
We compare PolSAM with ten state-of-the-art (SOTA) image segmentation methods across both datasets. First, we evaluate PolSAM using only pseudo-colored images (denoted as PolSAM w/o MVD) and compare it with five SAM-based methods, including those using optical Ke et al. (2024); Chai et al. (2023); Zhang et al. (2023a, e) and remote sensing images Zhang et al. (2024a), since the original SAM accepts only single inputs. Additionally, we compare PolSAM with five multimodal fusion algorithms Reza et al. (2024); He et al. (2023); Zhou et al. (2021); Zhang et al. (2023c); Zhou et al. (2023), as our method uses two inputs. The results are presented in Tables 1 and 2. The following is a detailed description of the results of the experiment.
Evaluation Metrics. For the PhySAR-Seg-1 dataset (Table 1), PolSAM outperforms other advanced SAM-based methods for both natural and remote sensing images. Specifically, PolSAM (w/o MVD) surpasses the top-performing method, RSAM-Seg Zhang et al. (2024a), with improvements of 1.44% in mAcc, 1.7% in mF1 score, and 2.09% in mIoU. Additionally, PolSAM shows significant superiority over other multimodal fusion methods, achieving SOTA performance.
For the PhySAR-Seg-2 dataset (Table 2), the SAM-LST method Chai et al. (2023) achieves the best performance among SAM-based approaches, with an mIoU of 42.82%. In comparison, PolSAM (w/o MVD) achieves 45.30%, improving by 2.48%. Due to the distinct nature of SAR data, existing SAM-based models struggle with optimal performance. PolSAM also outperforms multimodal fusion methods, achieving 0.48% higher mF1 and 2.38% better mIoU than GMNet Zhou et al. (2021), showcasing its effectiveness with diverse data inputs.
Additionally, the metrics for each category of the two datasets are visualized in Fig. 9, providing a clear comparison of the results. The blue bars represent our PolSAM(w/o MVD) method, corresponding to the blue-themed SAM-based approaches, while the red bars indicate our PolSAM(w/ MVD) method, aligned with the red-themed multimodal methods. Our method demonstrates three key strengths. First, both PolSAM (w/ MVD) and PolSAM (w/o MVD) outperform other methods across most categories. Second, PolSAM (w/ MVD) consistently surpasses PolSAM (w/o MVD), demonstrating the benefits of incorporating MVD. Third, PolSAM achieves optimal results across nearly all categories for the PhySAR-Seg-1 dataset. For the PhySAR-Seg-2 dataset, it shows significant improvement, particularly in the urban class where double-bounce scattering dominates despite small class proportion. This demonstrates that leveraging MVD in challenging scenarios substantially enhances segmentation accuracy.
Visualization. Fig. 10 presents segmentation results for the PhySAR-Seg-1 and PhySAR-Seg-2 datasets (first three and last three rows respectively), comparing SAM-based and multimodal fusion methods. PolSAM achieves the most refined segmentation results. MVD images highlight distinct terrain properties and high-level semantic information, outperforming pseudo-colored images in this regard. However, pseudo-colored images provide richer texture details. The multi-input complementarity at feature and semantic levels is clearly demonstrated. The proposed PolSAM explicitly leverages semantic information through fusion prompts, which guide the decoder toward superior segmentation performance.